### v2

@@ 3 Existence theorem of the UV limit measure
As a next step, we show that $Z_{\eta_{n}}(j)\to Z(j)$ as $n\to\infty$, whenever $\eta_{n}\in\mathcal{R}\subset\mathcal{S}$ ($n\in\mathbb{N}$) is an approximate identity in $\mathcal{S}$, and $j\in\mathcal{S}$. Indeed, for an approximate identity $(\eta_{n})_{n\in\mathbb{N}}$ one has $\eta^{t}_{n}\star j\overset{\mathcal{S}}{\to}j$ as $n\to\infty$, therefore
@@
    Staying on flat spacetime, but allowing strictly bandlimited regulators $\mathcal{R}=\mathcal{R}_{\mathcal{D}}$ only, the following can be said. Introduce the dense linear subspace $\mathcal{S}_{\mathcal{D}}$ of $\mathcal{S}$ consinsting of such elements whose Fourier transform reside in $\mathcal{D}$. Assign a vector topology to $\mathcal{S}_{\mathcal{D}}$ which is just the $\mathcal{D}$ topology on the Fourier transforms. Then, clearly $\mathcal{S}_{\mathcal{D}}\cong\mathcal{D}$ (the identificatiom map being the Fourier transformation), and thus $\mathcal{S}_{\mathcal{D}}$ is barrelled nuclear and therefore Bochner–Minlos theorem applies. Its continuous dual space $\mathcal{S}_{\mathcal{D}}^{\prime}$ equipped with the strong dual topology will contain $\mathcal{S}^{\prime}$ as a dense subspace, and $\mathcal{S}_{\mathcal{D}}^{\prime}\cong\mathcal{D}^{\prime}$. The analogy of Lemma 8 and Lemma 10 holds: by the construction of $\mathcal{D}$, for all $j\in\mathcal{S}_{\mathcal{D}}$ there exists some $\eta\in\mathcal{R}_{\mathcal{D}}$ and $\ell\in\mathcal{S}_{\mathcal{D}}$ such that $j=\eta\star\ell$, moreover, for all convergent sequences $j_{n}\in\mathcal{S}_{\mathcal{D}}$ ($n\in\mathbb{N}$) there exists some $\eta\in\mathcal{R}_{\mathcal{D}}$ and a convergent sequence $\ell_{n}\in\mathcal{S}_{\mathcal{D}}$ ($n\in\mathbb{N}$) such that $j_{n}=\eta\star\ell_{n}$ ($\forall\,n\in\mathbb{N}$). Consequently, the scheme of the proof of Theorem 14 can be copied when replacing $\mathcal{S}$ by $\mathcal{S}_{\mathcal{D}}$ and $\mathcal{R}_{\mathcal{S}}$ by $\mathcal{R}_{\mathcal{D}}$: it follows that there exists a corresponding positive definite sequentially $\mathcal{S}_{\mathcal{D}}$-continuous function $Z:\,\mathcal{S}_{\mathcal{D}}\to\mathbb{C}$ in the UV limit. If all the regularized measures $\mu_{\eta}$ ($\eta\in\mathcal{R}$) had second moments (which is always the case in QFT applications), using the RG equation, the estimate [33] (5.5.3) and the quadratic bound $1-\cos t\leq t^{2}$, by [16] Lemma21 it follows that $Z$ is also topologically $\mathcal{S}_{\mathcal{D}}$-continuous, and thus Bochner–Minlos theorem can be invoked. In summary, with $\mathcal{R}=\mathcal{R}_{\mathcal{D}}$, the UV limit measure $\mu$ of a Wilsonian RG flow still exists, but as a measure on $\mathcal{S}_{\mathcal{D}}^{\prime}$, having second moment. If this second moment distribution is tempered (i.e., does not blow up in frequency space faster than all polynomials), then $Z$ is $\mathcal{S}$-continuous, i.e. $\mu$ will be supported on $\mathcal{S}^{\prime}$.
@@ 4 Existence of UV limit of relative interaction potential
In QFT applications, one often needs to address the problematics of comparing Wilsonian RG flows of two related models. For instance, the flow of an interacting model, with respect to a reference flow corresponding to the Gaussian measure subordinate to the free Klein–Gordon operator $(-\Delta+m^{2})$, as outlined in Section 1. In particular, let $\mu_{\eta}$ and $\gamma_{\eta}$ be non-negative valued finite sigma-additive measures on $C_{\eta}[\mathcal{S}^{\prime}]$ ($\eta\in\mathcal{R}$), and assume that they form two Wilsonian RG flows in terms of Definition 2. Assume that the evolution of these two flows “cross” each-other in the sense that for a certain $\eta\in\mathcal{R}$ there exists some relative field renormalization factor $z_{\eta}\in\mathbb{R}^{+}$ between them, such that the measure $\mu_{\eta}$ is absolutely continuous to the reference measure $\tilde{\gamma}_{\eta}$ at this particular $\eta$, where $\tilde{\gamma}_{\eta}:=(z_{\eta})_{*}\,\gamma_{\eta}$ denotes the reference measure $\gamma_{\eta}$ re-expressed on the rescaled fields by the factor $z_{\eta}$. The absolute continuity, spelled out explicitely, means that there exists some Borel measurable function $f_{\eta}:\,C_{\eta}[\mathcal{S}^{\prime}]\to\overline{\mathbb{R}}_{0}^{+}$, such that $\mu_{\eta}=f_{\eta}\cdot\tilde{\gamma}_{\eta}$ holds. (In QFT, often this relative density function is expressed through its negative logarithm $V_{\eta}$, i.e. via $f_{\eta}=\mathrm{e}^{-V_{\eta}}$, called relative interaction potential.) Spelling out the ansatz $\mu_{\eta}=f_{\eta}\cdot\tilde{\gamma}_{\eta}$ means that for each Borel set $A$ of $C_{\eta}[\mathcal{S}^{\prime}]$ one has
$\mu_{\eta}(A)=\int\limits_{\varphi\in A}f_{\eta}(\varphi)\,\mathrm{d}\tilde{\gamma}_{\eta}(\varphi)=\int\limits_{\phi\in\overset{-1}{(z_{\eta}\,C_{\eta})}(A)}f_{\eta}(z_{\eta}\,C_{\eta}\phi)\,\mathrm{d}\gamma(\phi)$
@@
where at the last equality we used the fact that the family of reference measures $(\gamma_{\eta})_{\eta\in\mathcal{R}}$ form a Wilsonian RG flow (with UV limit $\gamma$ guaranteed by Corollary 15). The typical task is to try to characterize the evolution and UV limit of the flow $(\mu_{\eta})_{\eta\in\mathcal{R}}$ in terms of the reference flow $(\gamma_{\eta})_{\eta\in\mathcal{R}}$. [^3] As outlined at the end of Section 1, also formalized in Definition 2, without loss of generality, the running field renormalization factor can be merged into the flow of measures in a Wilsonian RG flow. That is, one can adapt a normalization convention in which it does not appear explicitely in the notation, when considering only a single flow. In this convention, it re-appears as a running relative field renormalization factor, when comparing two different flows against each-other. The answer to this question on physics ground is not evident at all, since the whole concept of Wilsonian RG flow was invented in order to overcome the difficulty of directly specifying a measure $\mu:=f\cdot\gamma$ in the UV limit, where the density $f$ would be defined by $f:=\mathrm{e}^{-V}$, with $V$ being some pointwise functional of distributional fields, such as $V(\phi)=g\,\int\phi^{4}$ according to the tentative definition. The development of the concept of Wilsonian RG flows was motivated by the expectation that such UV limit interaction potential $V$ may not exist in general, but only regularized interaction potentials $V_{\eta}$ should exist on the space of UV regularized fields $C_{\eta}[\mathcal{S}^{\prime}]$ ($\eta\in\mathcal{R}$). Our results will show that actually, the UV limit relative interaction potential $V$ exists, in a generalized sense.
@@
That is, $\mu$ is absolutely continuous with respect to the measure $(z_{\eta})_{*}\gamma$ with correspoding density function $f_{\eta}\circ C_{\eta}$.
@@
Assume that the conditions of Lemma 18 hold at some particular $\eta\in\mathcal{R}_{\mathcal{S}+}$. Then, it holds for all $\eta\in\mathcal{R}$, with $z_{\eta}=const$. Furthermore, in the normalization convention $z_{\eta}=1$ by means of Remark 19, it follows that $\mu$ is absolutely continuous to $\gamma$. That is, it follows that for all $\eta\in\mathcal{R}$ there exists a Borel measurable function $f_{\eta}:\,C_{\eta}[\mathcal{S}^{\prime}]\to\overline{\mathbb{R}}_{0}^{+}$, and an $f:\,\mathcal{S}^{\prime}\to\overline{\mathbb{R}}_{0}^{+}$, such that $\mu_{\eta}=f_{\eta}\cdot\gamma_{\eta}$ and $\mu=f\cdot\gamma$ holds. (In QFT, often $f_{\eta}$ and $f$ are characterized by their negative logarithm $V_{\eta}$ and $V$, respectively, i.e. $f_{\eta}=\mathrm{e}^{-V_{\eta}}$ and $f=\mathrm{e}^{-V}$, which are called interaction potentials.)
@@
By Lemma 18, assuming the normalization convention of Remark 19, it follows that the UV limit measure $\mu$ is absolutely continuous with respect to $\gamma$. That, by Radon–Nikodym theorem ([34] Theorem 6.10), is equivalent to the statement that for all Borel sets $B$ of $\mathcal{S}^{\prime}$ one has $\Big{(}\;\gamma(B)=0\;\Rightarrow\;\mu(B)=0\;\Big{)}$}. In order to prove the absolute continuity of $\mu_{\eta}$ with respect to $\gamma_{\eta}$ at arbitrary $\eta\in\mathcal{R}$, we show that for all Borel sets $B_{\eta}$ of $C_{\eta}[\mathcal{S}^{\prime}]$ one has $\Big{(}\;\gamma_{\eta}(B_{\eta})=0\;\Rightarrow\;\mu_{\eta}(B_{\eta})=0\;\Big{)}$}. For let $B_{\eta}$ be a Borel set of $C_{\eta}[\mathcal{S}^{\prime}]$ such that $\gamma_{\eta}(B_{\eta})=0$ holds. This, by means of Corollary 15 is equivalent to $\gamma(\overset{-1}{C_{\eta}}(B_{\eta}))=0$. Using the absolute continuity of $\mu$ with respect to $\gamma$, just derived above, by Radon–Nikodym theorem it follows that $\mu(\overset{-1}{C_{\eta}}(B_{\eta}))=0$. Applying Corollary 15, it follows then that $\mu_{\eta}(B_{\eta})=0$ holds, which completes the proof. ∎
@@
Let $\mathcal{R}=\mathcal{S}$ or $\mathcal{R}_{\mathcal{S}}$ or $\mathcal{R}_{\mathcal{S}+}$, and let $\mu_{\eta}$ and $\gamma_{\eta}$ be Wilsonian RG flows of non-negative valued finite sigma-additive measures on $C_{\eta}[\mathcal{S}^{\prime}]$ ($\eta\in\mathcal{R}$), with corresponding UV limit measures $\mu$ and $\gamma$ guaranteed by Corollary 15. If for some $\eta\in\mathcal{R}_{\mathcal{S}+}$ there exists $z_{\eta}\in\mathbb{R}^{+}$ and a Borel measurable function $f_{\eta}:\,C_{\eta}[\mathcal{S}^{\prime}]\to\overline{\mathbb{R}}_{0}^{+}$ such that $\mu_{\eta}:=f_{\eta}\cdot\tilde{\gamma}_{\eta}$ holds with $\tilde{\gamma}_{\eta}:=(z_{\eta})_{*}\,\gamma_{\eta}$, then this holds for all $\eta\in\mathcal{R}$ with a suitable $z_{\eta}$ and $f_{\eta}$, moreover $z_{\eta}$ can be chosen to be constant as a function of $\eta$. In the normalization convention of Remark 19, i.e. when $z_{\eta}=1$ by choice, there exists a Borel measurable function $f:\,\mathcal{S}^{\prime}\to\overline{\mathbb{R}}_{0}^{+}$ such that $\mu=f\cdot\gamma$. (In the QFT notation, $f_{\eta}=\mathrm{e}^{-V_{\eta}}$ and $f=\mathrm{e}^{-V}$ is used, and the pertinent $f$ and $V$ are the UV limit density function and interaction potential, respectively.) Furthermore, the following equivalent relations hold:
@@
    Let $\mathcal{R}=\mathcal{S}$ or $\mathcal{R}_{\mathcal{S}}$ or $\mathcal{R}_{\mathcal{S}+}$ and assume that at each $\eta\in\mathcal{R}$ for the studied Wilsonian RG flow the ansatz $\mu_{\eta}=\mathrm{e}^{-V_{\eta}}\cdot\gamma_{\eta}$ holds, where the reference measure $\gamma_{\eta}$ is the pushforward by $C_{\eta}$ of the canonical Gaussian measure on $\mathcal{S}^{\prime}$, associated to the Klein–Gordon operator $(-d^{2}\Delta+m^{2})$}, as outlined in Section 1 ($d,m\in\mathbb{R}^{+}$). One may allow in addition the parameters $d$ and $m$ to be possibly $\eta$-dependent. Then, due to Remark 37 (vi) and (viii), $d_{\eta}=const$ and $m_{\eta}=const$ must hold. That is, modulo the smoothing by $\eta\in\mathcal{R}$, a reference Gaussian subordinate to a Wilsonian RG flow cannot run.
@@
    Let $\mathcal{R}=\mathcal{R}_{\mathcal{D}}$, i.e. assume that only strictly bandlimited regulators are allowed. As shown in Remark 17 (ii), if $(\mu_{\eta})_{\eta\in\mathcal{R}}$ is a Wilsonian RG flow under this condition, its UV limit still extists, but as a measure $\mu$ on the space $\mathcal{S}_{\mathcal{D}}^{\prime}$. Assume that such flow satisfies the above ansatz in (i). Then, from Remark 37 (vi) and (viii) it still follows that $d_{\eta}=const$ and $m_{\eta}=const$ must hold. That is, the reference Gaussian subordinate to a Wilsonian RG flow, modulo smoothing, cannot run even when $\mathcal{R}=\mathcal{R}_{\mathcal{D}}$. It is an open question whether without further assumptions $\mu$ has support on $\mathcal{S}^{\prime}\subset\mathcal{S}_{\mathcal{D}}^{\prime}$ under the ansatz in (i), and if so, whether $\mu$ is absolutely continuous to the reference UV Gaussian $\gamma$. We intend to investigate these questions in a paper in preparation, focusing entirely on the bandlimited regulator case.

### v3

@@ 3 Existence theorem of the UV limit measure
[…]
@@
    Staying on flat spacetime, but allowing strictly bandlimited regulators $\mathcal{R}=\mathcal{R}_{\mathcal{D}}$ only, the following can be said. Introduce the dense linear subspace $\mathcal{S}_{\mathcal{D}}$ of $\mathcal{S}$ consisting of such elements whose Fourier transform reside in $\mathcal{D}$. Assign a vector topology to $\mathcal{S}_{\mathcal{D}}$ which is just the $\mathcal{D}$ topology on the Fourier transforms. Then, clearly $\mathcal{S}_{\mathcal{D}}\cong\mathcal{D}$ (the identification map being the Fourier transformation), and thus $\mathcal{S}_{\mathcal{D}}$ is barrelled nuclear and therefore Bochner–Minlos theorem applies. Its continuous dual space $\mathcal{S}_{\mathcal{D}}^{\prime}$ equipped with the strong dual topology will contain $\mathcal{S}^{\prime}$ as a dense subspace, and $\mathcal{S}_{\mathcal{D}}^{\prime}\cong\mathcal{D}^{\prime}$. The analogy of Lemma 8 and Lemma 10 holds: by the construction of $\mathcal{D}$, for all $j\in\mathcal{S}_{\mathcal{D}}$ there exists some $\eta\in\mathcal{R}_{\mathcal{D}}$ and $\ell\in\mathcal{S}_{\mathcal{D}}$ such that $j=\eta\star\ell$, moreover, for all convergent sequences $j_{n}\in\mathcal{S}_{\mathcal{D}}$ ($n\in\mathbb{N}$) there exists some $\eta\in\mathcal{R}_{\mathcal{D}}$ and a convergent sequence $\ell_{n}\in\mathcal{S}_{\mathcal{D}}$ ($n\in\mathbb{N}$) such that $j_{n}=\eta\star\ell_{n}$ ($\forall\,n\in\mathbb{N}$). Consequently, the scheme of the proof of Theorem 14 can be copied when replacing $\mathcal{S}$ by $\mathcal{S}_{\mathcal{D}}$ and $\mathcal{R}_{\mathcal{S}}$ by $\mathcal{R}_{\mathcal{D}}$: it follows that there exists a corresponding positive definite sequentially $\mathcal{S}_{\mathcal{D}}$-continuous function $Z:\,\mathcal{S}_{\mathcal{D}}\to\mathbb{C}$ in the UV limit. If all the regularized measures $\mu_{\eta}$ ($\eta\in\mathcal{R}$) had second moments (which is always the case in QFT applications), using the RG equation, the estimate [37] (5.5.3) and the quadratic bound $1-\cos t\leq t^{2}$, by [16] Lemma21 it follows that $Z$ is also topologically $\mathcal{S}_{\mathcal{D}}$-continuous, and thus Bochner–Minlos theorem can be invoked. In summary, with $\mathcal{R}=\mathcal{R}_{\mathcal{D}}$, the UV limit measure $\mu$ of a Wilsonian RG flow still exists, but as a measure on $\mathcal{S}_{\mathcal{D}}^{\prime}$, having second moment. If this second moment distribution is tempered (i.e., does not blow up in frequency space faster than all polynomials), then $Z$ is $\mathcal{S}$-continuous, i.e. $\mu$ will be supported on $\mathcal{S}^{\prime}$.
@@
Since the paper deals with Euclidean QFT models, a note on reflection positivity property is in order. Denote by $E_{\varepsilon+}:=\left\{x\in\mathbb{R}^{N}\,\big|\,x_{1}>\varepsilon\right\}$ the open half-spacetime with margin $\varepsilon\geq 0$. Reflection positivity of a probability measure $\mu$ on $\mathcal{S}^{\prime}$ would mean that for its Fourier transform $Z$ the following property holds: for all $j_{1},\dots,j_{m}\in\mathcal{D}$ with their support in $E_{0+}$, the matrix $\big(Z(j_{k}-\theta\,j_{l})\big)_{1\leq k,l\leq m}$ is positive semidefinite, where $\theta$ denotes the reflection of the $x_{1}$ spacetime coordinate (see e.g. [1] Section 6.1 (i)). The first question is: assuming the reflection positivity of the UV limit $\mu$ of a Wilsonian RG flow as in Corollary 15, what does it imply for the regularized measures $(\mu_{\eta})_{\eta\in\mathcal{R}}$ in the flow. Let $\mathcal{R}=\mathcal{S}$ and take an $\eta\in\mathcal{R}$ such that it is symmetric, non-negative valued, compactly supported, with its support contained in the unit ball of $\mathbb{R}^{N}$, and having unit integral. Its compressed version $(\eta_{\Lambda})_{1\leq\Lambda<\infty}$ in terms of Remark 11 (i) will define a $\mathcal{D}\to\mathcal{D}$ approximate identity, which will also be an $\mathcal{S}\to\mathcal{S}$ approximate identity (see e.g. [38] p.326). It is easily read off that $\mu_{\eta_{\Lambda}}$ will be reflection positive with margin $1/\Lambda$: for all $j_{1},\dots,j_{m}\in\mathcal{D}$ with their support in $E_{1/\Lambda\,+}$, the matrix $\big(Z_{\eta_{\Lambda}}(j_{k}-\theta\,j_{l})\big)_{1\leq k,l\leq m}$ is positive semidefinite. This follows simply from the identity
$\mathrm{supp}(\eta_{\Lambda}^{t}\star j)\subset\mathrm{supp}(\eta_{\Lambda}^{t})+\mathrm{supp}(j),$
the symmetricity of $\eta_{\Lambda}$, and (15). Conversely, if there is an approximate identity of the above kind, such that each $\mu_{\eta_{\Lambda}}$ is reflection positive with margin $1/\Lambda$, then $\mu$ will be reflection positive, by the weak convergence of $\mu_{\eta_{\Lambda}}$ to $\mu$. Similarly, if in addition, the regulator $\eta$ was chosen to be $O(N)$ invariant, the Euclidean invariance is inherited between the subflow $(\mu_{\eta_{\Lambda}})_{1\leq\Lambda<\infty}$ and the UV limit $\mu$, by analogous arguments.
@@ 4 Existence of UV limit of relative interaction potential
In QFT applications, one often needs to address the problematics of comparing Wilsonian RG flows of two related models. For instance, the flow of an interacting model, with respect to a reference flow corresponding to the Gaussian measure subordinate to the free Klein–Gordon operator $(-\Delta+m^{2})$, as outlined in Section 1. In particular, let $\mu_{\eta}$ and $\gamma_{\eta}$ be non-negative valued finite sigma-additive measures on $C_{\eta}[\mathcal{S}^{\prime}]$ ($\eta\in\mathcal{R}$), and assume that they form two Wilsonian RG flows in terms of Definition 2. Assume that the evolution of these two flows “cross” each-other in the sense that for a certain $\eta\in\mathcal{R}$ there exists some relative field renormalization factor $z_{\eta}\in\mathbb{R}^{+}$ between them, such that the measure $\mu_{\eta}$ is absolutely continuous to the reference measure $\tilde{\gamma}_{\eta}$ at this particular $\eta$, where $\tilde{\gamma}_{\eta}:=(z_{\eta})_{*}\,\gamma_{\eta}$ denotes the reference measure $\gamma_{\eta}$ re-expressed on the rescaled fields by the factor $z_{\eta}$. The absolute continuity, spelled out explicitly, means that there exists some Borel measurable function $f_{\eta}:\,C_{\eta}[\mathcal{S}^{\prime}]\to\overline{\mathbb{R}}_{0}^{+}$, such that $\mu_{\eta}=f_{\eta}\cdot\tilde{\gamma}_{\eta}$ holds. (In QFT, often this relative density function is expressed through its negative logarithm $V_{\eta}$, i.e. via $f_{\eta}=\rme^{-V_{\eta}}$, called relative interaction potential.) Spelling out the ansatz $\mu_{\eta}=f_{\eta}\cdot\tilde{\gamma}_{\eta}$ means that for each Borel set $A$ of $C_{\eta}[\mathcal{S}^{\prime}]$ one has
[…]
@@
where at the last equality we used the fact that the family of reference measures $(\gamma_{\eta})_{\eta\in\mathcal{R}}$ form a Wilsonian RG flow (with UV limit $\gamma$ guaranteed by Corollary 15). The typical task is to try to characterize the evolution and UV limit of the flow $(\mu_{\eta})_{\eta\in\mathcal{R}}$ in terms of the reference flow $(\gamma_{\eta})_{\eta\in\mathcal{R}}$. [^4] As outlined at the end of Section 1, also formalized in Definition 2, without loss of generality, the running field renormalization factor can be merged into the flow of measures in a Wilsonian RG flow. That is, one can adapt a normalization convention in which it does not appear explicitly in the notation, when considering only a single flow. In this convention, it re-appears as a running relative field renormalization factor, when comparing two different flows against each-other. The answer to this question on physics ground is not evident at all, since the whole concept of Wilsonian RG flow was invented in order to overcome the difficulty of directly specifying a measure $\mu:=f\cdot\gamma$ in the UV limit, where the density $f$ would be defined by $f:=\rme^{-V}$, with $V$ being some pointwise functional of distributional fields, such as $V(\phi)=g\,\int\phi^{4}$ according to the tentative definition. The development of the concept of Wilsonian RG flows was motivated by the expectation that such UV limit interaction potential $V$ may not exist in general, but only regularized interaction potentials $V_{\eta}$ should exist on the space of UV regularized fields $C_{\eta}[\mathcal{S}^{\prime}]$ ($\eta\in\mathcal{R}$). Our results will show that actually, the UV limit relative interaction potential $V$ exists, in a generalized sense.
@@
That is, $\mu$ is absolutely continuous with respect to the measure $(z_{\eta})_{*}\gamma$ with corresponding density function $f_{\eta}\circ C_{\eta}$.
@@
Assume that the conditions of Lemma 18 hold at some particular $\eta\in\mathcal{R}_{\mathcal{S}+}$. Then, it holds for all $\eta\in\mathcal{R}$, with $z_{\eta}=const$. Furthermore, in the normalization convention $z_{\eta}=1$ by means of Remark 19, it follows that $\mu$ is absolutely continuous to $\gamma$. That is, it follows that for all $\eta\in\mathcal{R}$ there exists a Borel measurable function $f_{\eta}:\,C_{\eta}[\mathcal{S}^{\prime}]\to\overline{\mathbb{R}}_{0}^{+}$, and an $f:\,\mathcal{S}^{\prime}\to\overline{\mathbb{R}}_{0}^{+}$, such that $\mu_{\eta}=f_{\eta}\cdot\gamma_{\eta}$ and $\mu=f\cdot\gamma$ holds. (In QFT, often $f_{\eta}$ and $f$ are characterized by their negative logarithm $V_{\eta}$ and $V$, respectively, i.e. $f_{\eta}=\rme^{-V_{\eta}}$ and $f=\rme^{-V}$, which are called interaction potentials.)
@@
[…]
@@
Let $\mathcal{R}=\mathcal{S}$ or $\mathcal{R}_{\mathcal{S}}$ or $\mathcal{R}_{\mathcal{S}+}$, and let $\mu_{\eta}$ and $\gamma_{\eta}$ be Wilsonian RG flows of non-negative valued finite sigma-additive measures on $C_{\eta}[\mathcal{S}^{\prime}]$ ($\eta\in\mathcal{R}$), with corresponding UV limit measures $\mu$ and $\gamma$ guaranteed by Corollary 15. If for some $\eta\in\mathcal{R}_{\mathcal{S}+}$ there exists $z_{\eta}\in\mathbb{R}^{+}$ and a Borel measurable function $f_{\eta}:\,C_{\eta}[\mathcal{S}^{\prime}]\to\overline{\mathbb{R}}_{0}^{+}$ such that $\mu_{\eta}:=f_{\eta}\cdot\tilde{\gamma}_{\eta}$ holds with $\tilde{\gamma}_{\eta}:=(z_{\eta})_{*}\,\gamma_{\eta}$, then this holds for all $\eta\in\mathcal{R}$ with a suitable $z_{\eta}$ and $f_{\eta}$, moreover $z_{\eta}$ can be chosen to be constant as a function of $\eta$. In the normalization convention of Remark 19, i.e. when $z_{\eta}=1$ by choice, there exists a Borel measurable function $f:\,\mathcal{S}^{\prime}\to\overline{\mathbb{R}}_{0}^{+}$ such that $\mu=f\cdot\gamma$. (In the QFT notation, $f_{\eta}=\rme^{-V_{\eta}}$ and $f=\rme^{-V}$ is used, and the pertinent $f$ and $V$ are the UV limit density function and interaction potential, respectively.) Furthermore, the following equivalent relations hold:
@@
    Let $\mathcal{R}=\mathcal{S}$ or $\mathcal{R}_{\mathcal{S}}$ or $\mathcal{R}_{\mathcal{S}+}$ and assume that at each $\eta\in\mathcal{R}$ for the studied Wilsonian RG flow the ansatz $\mu_{\eta}=\rme^{-V_{\eta}}\cdot\gamma_{\eta}$ holds, where the reference measure $\gamma_{\eta}$ is the pushforward by $C_{\eta}$ of the canonical Gaussian measure on $\mathcal{S}^{\prime}$, associated to the Klein–Gordon operator $(-d^{2}\Delta+m^{2})$}, as outlined in Section 1 ($d,m\in\mathbb{R}^{+}$). One may allow in addition the parameters $d$ and $m$ to be possibly $\eta$-dependent. Then, due to Remark 37 (vi) and (viii), $d_{\eta}=const$ and $m_{\eta}=const$ must hold. That is, modulo the smoothing by $\eta\in\mathcal{R}$, a reference Gaussian subordinate to a Wilsonian RG flow cannot run.
@@
    Let $\mathcal{R}=\mathcal{R}_{\mathcal{D}}$, i.e. assume that only strictly bandlimited regulators are allowed. As shown in Remark 17 (ii), if $(\mu_{\eta})_{\eta\in\mathcal{R}}$ is a Wilsonian RG flow under this condition, its UV limit still exists, but as a measure $\mu$ on the space $\mathcal{S}_{\mathcal{D}}^{\prime}$. Assume that such flow satisfies the above ansatz in (i). Then, from Remark 37 (vi) and (viii) it still follows that $d_{\eta}=const$ and $m_{\eta}=const$ must hold. That is, the reference Gaussian subordinate to a Wilsonian RG flow, modulo smoothing, cannot run even when $\mathcal{R}=\mathcal{R}_{\mathcal{D}}$. It is an open question whether without further assumptions $\mu$ has support on $\mathcal{S}^{\prime}\subset\mathcal{S}_{\mathcal{D}}^{\prime}$ under the ansatz in (i), and if so, whether $\mu$ is absolutely continuous to the reference UV Gaussian $\gamma$. We intend to investigate these questions in a paper in preparation, focusing entirely on the bandlimited regulator case.